# Lepton seesaw model in a modular $A_{4}$ symmetry

### v1

@@ Abstract
We propose a lepton seesaw model to get the same mass-origin of charged-leptons and neutrinos introducing a modular $A_{4}$ symmetry. In this scenario, the mass matrix of charged-leptons is induced via seesaw-like mechanism while the mass matrix of neutrino is realized via an inverse seesaw mechanism with help of vector-like fermions and an isospin singlet and triplet scalar field with nonzero vacuum expectation values. In analyzing the model we concentrate on fixed points of modulus $\tau$ that is favored by a flux compactification of Type IIB string theory. We search for allowed region at nearby these fixed points and find good solutions at nearby $\tau=\omega$ with several good predictions.
@@ I Introduction
Even though charged-leptons and the neutrinos are in the same component in the left-handed $SU(2)$ doublet of the standard model (SM), their masses are often discussed with different manners. This is because the masses of neutrinos are zero in the SM and we need to extend the SM so that we have (tiny but) non-zero neutrino masses. One of the simplest ways is to introduce heavy right-handed neutrinos that is analogy of the other three sectors, and the small neutrino masses can simply realized via seesaw mechanism Yanagida (1979); Minkowski (1977); Mohapatra and Senjanovic (1980); Zee (1980). But in this scenario, there is no relation between the mass matrix of neutrinos and charged-leptons. Thus it is interesting if we find a unified description of generating masses of them.
@@
In this paper, we study a possibility of the same origin of the charged-leptons and the neutrino mass matrices Lee et al. (2021), which is called ”lepton seesaw”, introducing several exotic fermions (vector-like fermions) and scalar fields (an isospin triplet scalar of hyper-charge 1 and singlet scalar of hyper-charge zero with nonzero vacuum expectations). In order to realize our model, we impose a modular flavor $A_{4}$ group that has a strong symmetry to get our desired terms as well as restricting some forms of mass matrices. Since 2017 Feruglio (2019), this symmetry, especially modular $A_{4}$, has widely been applied to various phenomenologies in flavor physics Criado and Feruglio (2018); Kobayashi et al. (2018); Okada and Tanimoto (2019); Nomura and Okada (2019); Okada and Tanimoto (2021a); de Anda et al. (2020); Novichkov et al. (2019); Nomura and Okada (2021); Okada and Orikasa (2019); Ding et al. (2019a); Nomura et al. (2020); Kobayashi et al. (2019); Asaka et al. (2020a); Zhang (2020); Ding et al. (2019b); Kobayashi et al. (2020); Nomura et al. (2021); Wang (2020); Okada and Shoji (2020); Okada and Tanimoto (2023); Behera et al. (2022a, b); Nomura and Okada (2022, 2020); Asaka et al. (2020b); Okada and Tanimoto (2021b); Nagao and Okada (2022); Okada and Tanimoto (2021c); Kang et al. (2022); Ding et al. (2024); Ding and King (2023); Nomura et al. (2024); Kobayashi et al. (2024); Petcov and Tanimoto (2024); Kobayashi and Tanimoto (2023); Nomura and Okada (2024); Qu and Ding (2024). In the model of this work, modular $A_{4}$ symmetry plays two important roles; (1) Forbidding unnecessary terms to realize lepton seesaw, (2) Providing predictions constraining structure of Yukawa couplings.
After constructing the model we perform numerical analysis to find some predictions in neutrino sector imposing experimental constraints. In our chi square numerical analysis, we concentrate on some specific region at nearby fixed points of modulus $\tau$ that is a new degree of freedom via modular symmetry. This fixed points are favored by a flux compactification of Type IIB string theory Ishiguro et al. (2021) and we find allowed region at nearby $\tau=\omega$ that reserves a remnant $Z_{3}$ symmetry, and obtain several predictions such as phases and masses on neutrino sector.
@@
This paper is organized as follows. In Sec. II, we review our setup of the modular $A_{4}$ assignments introducing new particles and explain how to realize our scenario by showing the charged-fermion sector and the neutral fermion sector. In Sec. III, we demonstrate chi-square numerical analyses concentrating on fixed points and show what kinds of predictions we obtain. Finally, we summarize and conclude in Sec. IV.
@@ II Model setup
$U(1)_{Y}$ $-\frac{1}{2}$ $1$ $\frac{1}{2}$ $-\frac{1}{2}$ $\frac{1}{2}$ $1$ 00
@@
$-k_{I}$ $+3$ $-2$ $-2$ $+2$ 00 $-4$ $-3$
Table 1: Charge assignments of the fermions and bosons under $SU(2)_{L}\otimes U(1)_{Y}\otimes A_{4}$ where $-k_{I}$ is the number of modular weight. Here, $\{\bm{1}\}=\{1,1^{\prime},1^{\prime\prime}\}$ indicates assignment of $A_{4}$ singlets.
Here, we review our model. At first, we introduce three isospin doublet vector-like leptons $L^{\prime}(\equiv[N,E]^{T}$) with $U(1)_{Y}=-1/2$, an isospin triplet scalar field $\Delta$ with $U(1)_{Y}=1$, and an isospin singlet scalar field $\varphi(\equiv(v^{\prime}+r)/\sqrt{2})$ with $U(1)_{Y}=0$, in addition to the SM fields. The SM Higgs is denoted by $H$ and its vacuum expectation value (VEV) is defined by $\langle H\rangle=[0,v/\sqrt{2}]^{T}$. We also require $\varphi$ and $\Delta$ to develop VEVs denoted by $v^{\prime}$ and $v_{\Delta}$ respectively. The SM right-handed leptons $\ell_{R}$, and $L^{\prime}$ are assigned to be $\{\bm{1}\}=\{1,1^{\prime},1^{\prime\prime}\}$ under the $A_{4}$ symmetry, while the left-handed SM leptons $L_{L}\equiv[\nu_{L},\ell_{L}]^{T}$ are assigned to be triplet under the $A_{4}$ symmetry. All the scalar fields are trivial singlet under the $A_{4}$ symmetry. In addition, we impose the modular weight for each field; $+3$ for $L_{L}$, $-2$ for $\overline{\ell_{R}}$, $-2$ for $\overline{L^{\prime}_{R}}$, $+2$ for ${L^{\prime}_{L}}$, 00 for $H$, $-4$ for $\Delta$, and $-3$ for $\varphi$. All fields and their assignments are summarized in TABLE 1. These assignments forbids undesired mass terms at tree-level such as $\overline{\ell_{R}}{\tilde{H}}L_{L}$, $\overline{L^{C}_{L}}(i\tau_{2})\Delta L_{L}$ and $\overline{L^{C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}$ that breaks our lepton seesaw scenario. These terms are forbidden as the following manner; $\overline{\ell_{R}}{\tilde{H}}L_{L}$ and $\overline{L^{C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}$ are not allowed since there is no weight $-1$ modular form; $\overline{L^{C}_{L}}(i\tau_{2})\Delta L_{L}$ is forbidden since there is no weight $-2$ modular form.
@@
$-{\cal L}_{\ell}=[\overline{\ell_{R}}{\tilde{H}}L^{\prime}_{L}]+[Y^{(2)}_{3}\overline{L^{\prime}_{R}}L_{L}\varphi]+[\overline{L^{\prime C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}]+[\overline{L^{\prime}_{R}}(i\tau_{2})\Delta^{\dagger}L^{\prime C}_{R}]+[M_{L^{\prime}}\overline{L^{\prime}_{R}}L^{\prime}_{L}]+{\rm h.c.},$ (1)
where $\tau_{2}$ is the second Pauli matrix, and []delimited-[][$\cdots$][ ⋯ ] represents a trivial $A_{4}$ singlet ${\bm{1}}$ abbreviating the free parameters that are explicitly shown later. $Y^{(2)}_{3}\equiv[y_{1},y_{2},y_{3}]^{T}$ is modular coupling that is uniquely fixed when modulus $\tau$ is determined Feruglio (2019). The Higgs potential is given by
${\cal V}$ $=-\mu_{H}^{2}|H|^{2}-\mu_{\Delta}^{2}{\rm Tr}[\Delta^{\dagger}\Delta]-\mu_{\varphi}|\varphi|^{2}+\mu(H^{T}(i\tau_{2})\Delta^{\dagger}H+{\rm h.c.})$
$+\lambda_{H}|H|^{4}+\lambda_{\Delta}{\rm Tr}[|\Delta^{\dagger}\Delta|^{2}]+\lambda^{\prime}_{\Delta}{\rm det}[\Delta^{\dagger}\Delta]+\lambda_{\varphi}|\varphi|^{4}+\lambda_{H\Delta}|H|^{2}{\rm Tr}[\Delta^{\dagger}\Delta]$
$+\lambda^{\prime}_{H\Delta}\sum_{i}^{3}(H^{\dagger}\tau_{i}H){\rm Tr}[\Delta^{\dagger}\tau^{i}\Delta]{+\lambda_{H\varphi}|H|^{2}|\varphi|^{2}}+\lambda_{\Delta\varphi}{\rm Tr}[\Delta^{\dagger}\Delta]|\varphi|^{2},$ (2)
where $\tau_{i}\ (i=1,2,3)$ are Pauli matrices and some of free parameters includes modular forms; e.g., $\mu_{\Delta}\equiv\mu^{\prime}_{\Delta}/(i\tau-i\tau^{*})^{4}$. The scalar potential is similar to that of Higgs triplet model with a little modification due to one scalar singlet extension. In this work we assume the parameters in the potential satisfy phenomenological constraints and realize VEVs of $H$, $\Delta$ and $\varphi$.
@@ II.1 Charged-lepton mass matrix
$[\overline{\ell_{R}}{\tilde{H}}L^{\prime}_{L}]+[Y^{(2)}_{3}\overline{L^{\prime}_{R}}L_{L}\varphi]+[M_{L^{\prime}}\overline{L^{\prime}_{R}}L^{\prime}_{L}]\supset m\overline{\ell_{R}}E_{L}+m^{\prime}\overline{E_{R}}\ell_{L}+M_{L^{\prime}}\overline{E^{\prime}_{R}}E_{L},$ (3)
@@
where $\{a_{e},b_{e},c_{e}\}$, $\{f_{1},f_{2},f_{3}\}$, and $\{M_{1},M_{2},M_{3}\}$ can be real parameters without loss of generality. Then, in basis of $[\ell_{L},E_{L}]^{T}$, the mass matrix of charged-lepton sector is given by
${\cal M}_{e}=\left(\begin{array}[]{cc}0&m\\ m^{\prime}&M_{L^{\prime}}\end{array}\right)$ (19)
When we assume to be $m,m^{\prime}\ll M_{L^{\prime}}$ among mass matrices, we can treat the charged-lepton mass matrix as a seesaw-like as follows. First of all, ${\cal M}_{e}$ is diagonalized by
@@
Due to the mass hierarchies among $m,m^{\prime},M_{L^{\prime}}$, we write the mixing matrix and mass matrices as follows:
$O_{L}\simeq\left(\begin{array}[]{cc}1&-\theta_{L}\\ \theta_{L}&1\end{array}\right),\quad\theta_{L}\approx-M^{-1}m^{\prime},$ (25)
@@
Since $M_{L^{\prime}}$ is diagonal, we do not need to diagonalize it more. $m_{e}$ is diagonalized by $D_{e}=V_{e_{R}}m_{e}V_{e_{L}}$. Therefore, $|D_{e}|^{2}=V^{\dagger}_{e_{L}}m^{\dagger}_{e}m_{e}V_{e_{L}}$ where $D_{e}=$ diag.$(m_{e},m_{\mu},m_{\tau})$. Since $m_{e}$ is given by

### v2

@@ Abstract
We propose a lepton seesaw model to get the same mass-origin of charged-leptons and neutrinos introducing a modular $A_{4}$ symmetry. In this scenario, the mass matrix of charged-leptons is induced via seesaw-like mechanism while the mass matrix of neutrino is realized via an inverse seesaw mechanism with help of vector-like fermions and an isospin singlet and triplet scalar fields with nonzero vacuum expectation values. In analyzing the model we concentrate on fixed points of modulus $\tau$ that is favored by a flux compactification of Type IIB string theory. We search for allowed region at nearby these fixed points and find good solutions at nearby $\tau=\omega$ with several good predictions.
@@ I Introduction
Even though charged-leptons and the neutrinos are in the same component in the left-handed $SU(2)$ doublet of the standard model (SM), their masses are often discussed with different manners. This is because the masses of neutrinos are zero in the SM and we need to extend the SM so that we have (tiny but) non-zero neutrino masses. One of the simplest ways is to introduce heavy right-handed neutrinos that is analogy of the other three sectors, and the small neutrino masses can simply be realized via seesaw mechanism Yanagida (1979); Minkowski (1977); Mohapatra and Senjanovic (1980); Zee (1980). But in this scenario, there is no relation between the mass matrix of neutrinos and charged-leptons. Thus it is interesting if we find a unified description of generating masses of them.
@@
In this paper, we study a possibility of the same origin for the charged-leptons and the neutrino mass matrices Lee et al. (2021), which is called ”lepton seesaw”, introducing several exotic fermions (vector-like fermions) and scalar fields (an isospin triplet scalar of hyper-charge 1 and singlet scalar of hyper-charge zero with nonzero vacuum expectation values). In order to realize our model, we impose a modular flavor $A_{4}$ group that has a promising symmetry to get our desired terms as well as restricting some forms of mass matrices. Since 2017 Feruglio (2019), this symmetry, especially modular $A_{4}$, has widely been applied to various phenomenology in flavor physics Criado and Feruglio (2018); Kobayashi et al. (2018); Okada and Tanimoto (2019); Nomura and Okada (2019); Okada and Tanimoto (2021c); de Anda et al. (2020); Novichkov et al. (2019); Nomura and Okada (2021); Okada and Orikasa (2019); Ding et al. (2019b); Nomura et al. (2020); Kobayashi et al. (2019); Asaka et al. (2020a); Zhang (2020); Ding et al. (2019a); Kobayashi et al. (2020); Nomura et al. (2021); Wang (2020); Okada and Shoji (2020); Okada and Tanimoto (2023); Behera et al. (2022b, a); Nomura and Okada (2022, 2020); Asaka et al. (2020b); Okada and Tanimoto (2021a); Nagao and Okada (2022); Okada and Tanimoto (2021b); Kang et al. (2022); Ding et al. (2024); Ding and King (2023); Nomura et al. (2024); Kobayashi et al. (2024); Petcov and Tanimoto (2024); Kobayashi and Tanimoto (2023); Nomura and Okada (2024); Qu and Ding (2024). Furthermore, a framework of non-holomorphic modular symmetry was recently developed Qu and Ding (2024) which helps to construct simpler models without supersymmetry. In the model of this work, the non-holomorphic modular $A_{4}$ symmetry plays two important roles; (1) Forbidding unnecessary terms to realize lepton seesaw, (2) Providing predictions constraining structure of Yukawa couplings. After constructing the model we perform numerical analysis to find some predictions in neutrino sector imposing experimental constraints. In our chi square numerical analysis, we concentrate on some specific regions at nearby fixed points of modulus $\tau$ that is a new degree of freedom via modular symmetry. These fixed points are favored by a flux compactification of Type IIB string theory Ishiguro et al. (2021). We find allowed regions at nearby $\tau=\omega$ that preserves a remnant $Z_{3}$ symmetry, and obtain several predictions such as phases and masses in neutrino sector.
@@
This paper is organized as follows. In Sec. II, we review our setup of the modular $A_{4}$ assignments introducing new particles and explain how to realize our scenario by showing the mass generation mechanisms for charged-fermion sector and the neutral fermion sector. In Sec. IV, we demonstrate chi-square numerical analyses concentrating on fixed points and show what kinds of predictions we obtain. Finally, we summarize and conclude in Sec. V.
@@ II Model setup
$U(1)_{Y}$ $\frac{1}{2}$ 11 $\frac{1}{2}$ $\frac{1}{2}$ $\frac{1}{2}$ 11 0
@@
$k_{I}$ 1/2{-1/2} 2-2 2-2 +2+2 0 0{0} 1/2{1/2}
Table 1: Charge assignments of the fermions and bosons under $SU(2)_{L}\otimes U(1)_{Y}\otimes A_{4}$ where $k_{I}$ is the number of modular weight. Here, $\{\bm{1}\}=\{1,1^{\prime},1^{\prime\prime}\}$ indicates assignment of $A_{4}$ singlets.
Here, we review our model based on the non-holomorphic modular $A_{4}$ symmetry. At first, we introduce three isospin doublet vector-like leptons $L^{\prime}(\equiv[N,E]^{T}$) with hypercharge Y=1/2Y=-1/2, an isospin triplet scalar field $\Delta$ with $Y=1$, and an isospin singlet scalar field $\varphi(\equiv(v^{\prime}+r)/\sqrt{2})$ with $U(1)_{Y}=0$, in addition to the SM fields. The SM Higgs is denoted by $H$ and its vacuum expectation value (VEV) is defined by $\langle H\rangle=[0,v/\sqrt{2}]^{T}$. We also require $\varphi$ and $\Delta$ to develop VEVs denoted by $v^{\prime}$ and $v_{\Delta}$ respectively. The SM right-handed leptons $\ell_{R}$, and $L^{\prime}$ are assigned to be $\{\bm{1}\}=\{1,1^{\prime},1^{\prime\prime}\}$ under the $A_{4}$ symmetry, while the left-handed SM leptons $L_{L}\equiv[\nu_{L},\ell_{L}]^{T}$ are assigned to be triplet under the $A_{4}$ symmetry.  [^1] Although we adopt the non-holomorphic modular $A_{4}$ symmetry in this work, it is possible to obtain the same structure using holomorphic case. When one needs to work on a supersymmetric theory, main discussions below are totally maintained by changing assignments for modular weight into +3+3 for $L_{L}$, $\Delta$ for 4-4, and 3-3 for $\varphi$. All the scalar fields are trivial singlet under the $A_{4}$ symmetry. In addition, we impose the modular weight for each field; 1/2-1/2 for $L_{L}$, 2-2 for $\overline{\ell_{R}}$, 2-2 for $\overline{L^{\prime}_{R}}$, +2+2 for $L^{\prime}_{L}$}, 0 for $H$, 0 for $\Delta$, and $1/2$ for $\varphi$. All fields and their charge assignments are summarized in TABLE 1. These assignments forbid undesired mass terms at tree-level such as $\overline{\ell_{R}}{\tilde{H}}L_{L}$, $\overline{L^{C}_{L}}(i\tau_{2})\Delta L_{L}$ and $\overline{L^{C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}$ due to non-even number of modular weights.
@@
$-{\cal L}_{\ell}=[\overline{\ell_{R}}{H^{\dagger}}L^{\prime}_{L}]+[Y^{(2)}_{3}\overline{L^{\prime}_{R}}L_{L}\varphi]+{[Y^{(-4)}_{1}\overline{L^{\prime C}_{L}}(i\tau_{2})\Delta L^{\prime}_{L}]}+{[Y^{(4)}_{1,1^{\prime}}\overline{L^{\prime}_{R}}(i\tau_{2})\Delta^{\dagger}L^{\prime C}_{R}]}+[M_{L^{\prime}}\overline{L^{\prime}_{R}}L^{\prime}_{L}]+{\rm h.c.},$ (1)
where $\tau_{2}$ is the second Pauli matrix, and $\cdots$] represents a trivial $A_{4}$ singlet $\bm{1}$} abbreviating the free parameters that are explicitly shown later. $Y^{(2)}_{3}\equiv[y_{1},y_{2},y_{3}]^{T}$ is modular coupling that is uniquely fixed when modulus $\tau$ is determined Feruglio (2019). The Higgs potential is given by
${\cal V}$ $=-\mu_{H}^{2}|H|^{2}+\mu_{\Delta}^{2}{\rm Tr}[\Delta^{\dagger}\Delta]-\mu_{\varphi}^{2}|\varphi|^{2}+\mu(H^{T}(i\tau_{2})\Delta^{\dagger}H+{\rm h.c.})$
$+\lambda_{H}|H|^{4}+\lambda_{\Delta}{\rm Tr}[|\Delta^{\dagger}\Delta|^{2}]+\lambda^{\prime}_{\Delta}{\rm det}[\Delta^{\dagger}\Delta]+\lambda_{\varphi}|\varphi|^{4}{+(\tilde{\lambda}_{\varphi}Y_{1}^{(-2)}\varphi^{4}+h.c.)}$
$+\lambda_{H\Delta}|H|^{2}{\rm Tr}[\Delta^{\dagger}\Delta]+\lambda^{\prime}_{H\Delta}\sum_{i}^{3}(H^{\dagger}\tau_{i}H){\rm Tr}[\Delta^{\dagger}\tau^{i}\Delta]{+\lambda_{H\varphi}|H|^{2}|\varphi|^{2}}+\lambda_{\Delta\varphi}{\rm Tr}[\Delta^{\dagger}\Delta]|\varphi|^{2},$ (2)
where $\tau_{i}\ (i=1,2,3)$ are Pauli matrices and some of free parameters include modular forms; e.g., $\mu_{\varphi}\equiv\mu^{\prime}_{\varphi}(i\tau-i\tau^{*})^{1/2}$. The scalar potential is similar to that of Higgs triplet model with a little modification due to one scalar singlet extension. The scalar fields can be written by
$H=\begin{pmatrix}G^{+}\\ \frac{1}{\sqrt{2}}(v+h+iG)\end{pmatrix},\quad\Delta=\begin{pmatrix}\frac{\delta^{+}}{\sqrt{2}}&\delta^{++}\\ \frac{1}{\sqrt{2}}(v_{\Delta}+\delta^{0}+i\eta)&-\frac{\delta^{+}}{\sqrt{2}}\end{pmatrix},\quad\varphi=\frac{1}{\sqrt{2}}(v^{\prime}+\phi+i\chi),$ (3)
where $G^{\pm}$ and $G$ are Nambu-Goldstone bosons absorbed by the massive SM gauge fields $W^{\pm}$ and $Z$. The VEVs of the scalar fields are obtained by solving the conditions
$\frac{\partial V}{\partial v}\Bigg|_{0}=\frac{\partial V}{\partial v^{\prime}}\Bigg|_{0}=\frac{\partial V}{\partial v_{\Delta}}\Bigg|_{0}=0,$ (4)
with $|_{0}$ indicating all the component fields to be zero. The stable vacuum can be obtained by assuming $\mu^{2}_{\{H,\Delta,\varphi\}}>0$, $\mu>0$, $\lambda_{\{H,\varphi\}}>0$ where we also assume $\{\tilde{\lambda}_{\varphi},\lambda_{H\varphi}\}\to 0$ for simplicity. The VEVs are then obtained such that
$v\simeq\sqrt{\frac{\mu^{2}_{H}}{\lambda_{H}}},\quad v^{\prime}\simeq\sqrt{\frac{\mu^{2}_{\varphi}}{\lambda_{\varphi}}},\quad v_{\Delta}\simeq\frac{\mu v^{2}}{\sqrt{2}\mu_{\Delta}^{2}}.$ (5)
The triplet VEV can be small taking $v^{2}\ll\mu^{2}_{\Delta}$ by type-II seesaw mechanism. The masses of scalar fields are also simply given by
$m_{h}\simeq\sqrt{\lambda_{H}}v,\quad m_{\{\phi,\chi\}}\simeq\sqrt{\lambda_{\varphi}}v^{\prime},\quad m_{\{\delta^{0},\eta,\delta^{+},\delta^{++}\}}\simeq\mu_{\Delta}.$ (6)
In our analysis below the extra scalar masses are not important and we just choose them to be larger than TeV scale in order to satisfy current collider constraints.
@@ II.1 Charged-lepton mass matrix
$[\overline{\ell_{R}}{H^{\dagger}}L^{\prime}_{L}]+[Y^{(2)}_{3}\overline{L^{\prime}_{R}}L_{L}\varphi]+[M_{L^{\prime}}\overline{L^{\prime}_{R}}L^{\prime}_{L}]$
$\equiv{a_{e}}\overline{e_{R}}{H^{\dagger}}L^{\prime}_{L_{1}}+{b_{e}}\overline{\mu_{R}}{H^{\dagger}}L^{\prime}_{L_{2}}+{c_{e}}\overline{\tau_{R}}{H^{\dagger}}L^{\prime}_{L_{3}}$
$+f_{1}\overline{L^{\prime}_{R_{1}}}[y_{1}L_{L_{1}}+y_{2}L_{L_{3}}+y_{3}L_{L_{2}}]\varphi{+f_{2}\overline{L^{\prime}_{R_{2}}}[y_{3}L_{L_{3}}+y_{1}L_{L_{2}}+y_{2}L_{L_{1}}]\varphi}$
${+f_{3}\overline{L^{\prime}_{R_{3}}}[y_{2}L_{L_{2}}+y_{1}L_{L_{3}}+y_{3}L_{L_{1}}]\varphi}+M_{1}\overline{L^{\prime}_{R_{1}}}L^{\prime}_{L_{1}}+M_{2}\overline{L^{\prime}_{R_{2}}}L^{\prime}_{L_{2}}+M_{3}\overline{L^{\prime}_{R_{3}}}L^{\prime}_{L_{3}}$ (7)
$\supset m\overline{\ell_{R}}E_{L}+m^{\prime}\overline{E_{R}}\ell_{L}+M_{L^{\prime}}\overline{E_{R}}E_{L},$ (8)
@@
where $\{a_{e},b_{e},c_{e}\}$, $\{f_{1},f_{2},f_{3}\}$, and $\{M_{1},M_{2},M_{3}\}$ can be real parameters without loss of generality. Then, in basis of $\ell_{L},E_{L}]^{T}$, the mass matrix of charged-lepton sector is given by
${\cal M}_{e}=\left(\begin{array}[]{cc}0&m\\ m^{\prime}&M_{L^{\prime}}\end{array}\right).$ (24)
When we assume relations $m,m^{\prime}\ll M_{L^{\prime}}$ among mass matrices,  [^2] It is reasonable assumption since $M_{L^{\prime}}$ is mass parameter in original Lagrangian and would be large. However these hierarchies cannot be controlled by any flavor symmetry and one might reply on the other mechanism such as radiatively induced scenarios, e.g. ref. Das et al. (2017). Here, it is beyond our scope. we can treat the charged-lepton mass matrix as a seesaw-like as follows. First of all, ${\cal M}_{e}$ is diagonalized by
@@
Due to the mass hierarchies among $m$, $m^{\prime}$ and $M_{L^{\prime}}$, we write the mixing matrix and mass matrices as follows:
$O_{L}\simeq\left(\begin{array}[]{cc}1&-\theta_{L}\\ \theta_{L}&1\end{array}\right),\quad\theta_{L}\approx-M_{L^{\prime}}^{-1}m^{\prime},$ (30)
@@
[…]